OpenStudy (anonymous):

Help? g(x) = −2x + 4, find g(−5)

OpenStudy (anonymous):

g(-5) = -2(-5) + 4 g(-5) = 10 + 4 g(-5) = 14 <-- Is the answer "14"??

OpenStudy (helder_edwin):

yes.